OpenStudy (anonymous):

Prove the identity 1+sinx/cosx+cosx/1+sinx=2/cosx

OpenStudy (anonymous):

|sinx^2 + cosx^2 = 1 1+ 2sinx + 1 = 2 + 2sinx 2 + 2sinx = 2 + 2sinx

OpenStudy (anonymous):

Sorry hershey we should prove that the left hand side is identical to the right hand side

OpenStudy (anonymous):

Get common denominators on the left then add The LCD is cos x(1 + sin x) so you get (1 + sin x)(1 + sin x) / [cos x (1 + sin x)] + cos x (cos x) / [cos x (1 + sin x)] = 1 + 2 sin x + sin^2 x + cos^2 x on top and since sin^2 + cos^2 = 1, this is 2 + 2 sin x or 2(1 + sin x) over [cos x (1 + sin x)] and you cancel the (1 + sin x) leaving what you want

OpenStudy (anonymous):

I can give you 1.Results 2.Solution 3.Input 4. Atl form (Alternate) 5. Atl form assuming x is real: 6. Number line Lol :D

OpenStudy (anonymous):

\[(1+sinx)/cosx + cosx/(1+sinx)=2/cosx\] \[cosx/(1+sinx)=2/cosx- (1+sinx)/cosx\] \[cosx/(1+sinx)=(2-1-sinx)/cosx\] \[\cos ^{2}x=(1+sinx)(1-sinx)\] \[\cos ^{2}x=1-\sin ^{2}x\] \[\sin ^{2}x+\cos ^{2}x=1\] \[1=1\] Therefore, LHS=RHS. Hope this helps.

OpenStudy (anonymous):

Well thanks a lot guys vengeance kisses method is right and correct boss we end up having2/cosx which is on the ryt hand side :)

OpenStudy (anonymous):

So as long you get the right the right answer, I don't care who gives it :)

OpenStudy (anonymous):

thats more like it :)